A contractor has purchased a wheel loader for $115,000 and plans to use it 2,000 hours per year. The cost of one set of tires is $25,000. At this usage rate, the contractor anticipates disposing of the loader after using it for 10 years and realizing a salvage value of $35,000. The flywheel horsepower rating of the loader’s diesel engine is 105 horsepower. The interest rate is 10%. The loader operator will earn $34.00 per hour including fringe benefits, and diesel fuel costs $1.20 per gallon.

How much is the hourly equipment repair cost if the annual cost of repairs equals 70% of straight-line depreciation?
Answer #1

Cost of wheel loader=$115,000

Salvage Value =$35,000

Useful Life in years=10

Annual Depreciation =(115000-35000)/10=$8,000

Annual Cost of Repair =70%*8000=$5,600

Annual usage in hours=2,000

Hourly Equipment Repair Cost =5600/2000=$2.80

Hourly Equipment Repair Cost $2.80